\begin{thebibliography} {99}

\bibitem{BCM} A. Banerjee, A. Char, B. Mondal, Spectra of general hypergraphs, Linear Algebra Appl. 518 (2017) 14-30.

\bibitem{BN} B. Bollob\'{a}s, V. Nikiforov, Cliques and the spectral radius, J. Combin. Theory Ser. B 97 (2007) 859-865.

\bibitem{RP} S.R. Bulo, M. Pelillo, New bounds on the clique number of graphs based on spectral hypergraph theory, learning and intelligent optimization (2009) 45-58.

\bibitem{DW} C.X. Duan, L.G. Wang, The $\alpha$-spectral radius of $f$-connected general hypergraphs, Appl. Math. Comput. 382 (2020) 125336.

\bibitem{DWL} C.X. Duan, L.G. Wang, X.H. Li, Some properties of the Signless Laplacian and normalized Laplacian tensors of general hypergraphs, Taiwanese J. Math. 24 (2) (2020) 265-281.

\bibitem{DWXL} C.X. Duan, L.G. Wang, P. Xiao, X.H. Li, The (signless Laplacian) spectral radius (of subgraphs) of uniform hypergraphs, Filomat 33 (15) (2019) 4733-4745.

\bibitem{FTP} Y.Z. Fan, Y.Y. Tan, X.X. Peng, A.H. Liu, Maximizing spectral radii of uniform hypergraphs with few edges, Discuss. Math. Graph Theory 36 (2016) 845¨C856.

\bibitem{HLP} G.H. Hardy, J.E. Littlewood, G. P\'{o}lya, Inequalities, second ed., Cambridge University Press, 1988.

\bibitem{HL} P. Hansen, C. Lucas, Bounds and conjectures for the signless Laplacian index of graphs, Linear Algebra Appl. 432 (2010) 3319-3336.

\bibitem{HJZ} B. He, Y.L. Jin, X.D. Zhang, Sharp bounds for the signless Laplacian spectral radius in terms of clique number, Linear Algebra Appl. 438 (2013) 3851-3861.

\bibitem{HCZ} Y. Hou, A. Chang, L. Zhang, A homogeneous polynomial associated with general hypergraphs and its applications, Linear Algebra Appl. 591 (2020) 72-86.

\bibitem{MS} T.S. Motzkin, E.G. Straus, Maxima for graphs and a new proof of a theorem of Tur\'{a}n, Canad. J. Math. 17 (1965) 533-540.

\bibitem{L} L.H. Lim, Singular values and eigenvalues of tensors: a variational approach, in: Proceedings of the IEEE International Workshop on Computational Advances in Multi-Sensor Adaptive Processing (CAMSAP 05) 1 (2005) 129-132.

\bibitem{LZM} H.Y. Lin, B. Zhou, B. Mo, Upper bounds for $H$- and $Z$-spectral radii of uniform hypergraphs, Linear Algebra Appl. 510 (2016) 205-221.

\bibitem{LL} J.P. Liu, B.L. Liu, The maximum clique and the signless Laplacian eigenvalues, Czech. Math. J. 58 (133) (2008) 1233-1240.

\bibitem{LLT} M. Lu, H.Q. Liu, F. Tian, Laplacian spectral bounds for clique and independence numbers of graphs, J. Combin. Theory Ser. B 97 (2007) 726-732.

\bibitem{N} V. Nikiforov, More spectral bounds on the clique and independence numbers, J. Combin. Theory Ser. B 99 (2009) 819-826.

\bibitem{Q1} L.Q. Qi, Eigenvalues of a real supersymmetric tensor, J. Symbolic Comput. 40 (2005) 1302-1324.

\bibitem{Q2} L.Q. Qi, $H^{+}$-eigenvalues of Laplacian and signless Laplacian tensors, Commun. Math. Sci. 12 (2014) 1045-1064.

\bibitem{W} H.S. Wilf, Spectral bounds for the clique and independence numbers of graphs, J. Combin. Theory Ser. B 40 (1986) 113-117.

\bibitem{XW1} P. Xiao, L.G. Wang, Y. Lu, The maximum spectral radii of uniform supertrees with given degree sequences, Linear Algebra Appl. 523 (2017) 33-45.

\bibitem{XW2} P. Xiao, L.G. Wang, The effect on the (signless Laplacian) spectral radii of uniform hypergraphs by subdividing an edge, Discrete Appl. Math. 283 (2020) 444-455.

\bibitem{XQ} J.S. Xie, L.Q. Qi, The clique and coclique numbers' bounds based on the $H$-eigenvalues of uniform hypergraphs, Int. J. Numer. Anal. Model. 12 (2015) 318-327.

\bibitem{ZLKB} W. Zhang, L.L. Liu, L.Y. Kang, Y.Q. Bai, Some properties of the spectral radius for general hypergraphs, Linear Algebra Appl. 513 (2017) 103-119.


\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{enumerate}
\usepackage{amssymb,a4wide,latexsym,makeidx,epsfig,fleqn}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{enumerate}
\usepackage{float}
\usepackage{color}
\usepackage{amsmath}
\allowdisplaybreaks[4]
\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}
\newtheorem{lemma}[theorem]{Lemma}

\begin{document}
\textwidth 150mm \textheight 225mm
\title{The bounds of the spectral radius of general hypergraphs in terms of clique number
\thanks{Supported by the National Natural Science Foundation of China (No. 11871398) and the Seed Foundation of Innovation and Creation for Graduate Students in Northwestern Polytechnical University (No. CX2020190).}}
\author{{Cunxiang Duan$^{a,b}$, Ligong Wang$^{a,b,}$\footnote{Corresponding author.} }\\
{\small $^{a}$School of Mathematics and Statistics, Northwestern
Polytechnical University,} \\ {\small Xi'an, Shaanxi 710129, P.R. China.} \\ {\small $^{b}$Xi'an-Budapest Joint Research Center for Combinatorics, Northwestern Polytechnical University,}
\\ {\small Xi'an, Shaanxi 710129, P.R. China.}
 \\{\small E-mail: cxduanmath@163.com; lgwangmath@163.com}\\}
\date{}
\maketitle
\begin{center}
\begin{minipage}{120mm}
\vskip 0.3cm
\begin{center}
{\small {\bf Abstract}}
\end{center}
{\small The spectral radius (or the signless Laplacian spectral radius) of a general hypergraph is the maximum modulus of the eigenvalues of its adjacency (or  its signless Laplacian) tensor. In this paper, we firstly obtain a lower bound of the spectral radius (or the signless Laplacian spectral radius) of general hypergraphs in terms of clique number. Moreover, we present a relation between a homogeneous polynomial and the clique number of general hypergraphs. As an application, we finally obtain an upper bound of the spectral radius of general hypergraphs in terms of clique number.

\vskip 0.1in \noindent {\bf Key Words}: \  spectral radius, clique number, general hypergraphs \vskip
0.1in \noindent {\bf AMS Subject Classification (2020)}: \  15A42, 05C50.}
\end{minipage}
\end{center}

\section{Introduction}
Let $G$ be a general hypergraph with vertex set $V=V(G)=\left\{v_{1}, v_{2}, \ldots, v_{n}\right\}$ and edge set $E=E(G)=\left\{e_{1}, e_{2}, \dots, e_{m}\right\},$ where $e_{1}, e_{2},\ldots, e_{m}\subseteq V(G).$ If $V'\subseteq V$ and $E'\subseteq E$, then $H = (V', E')$ is called a subhypergraph of $G.$ The rank (resp., co-rank) of $G$ is $rank(G)=\max \{|e|\colon\ e \in E\}$ (resp., $corank(G)=\min\{|e|\colon\ e \in E\}$).  If the rank and the co-rank of a hypergraph $G$ are equal to $k$, then $G$ is called $k$-uniform. Obviously, a graph is a $2$-uniform hypergraph. For two vertices $u, v \in V,$ if there exists an edge $e$ such that $u, v \in e \in E,$ then vertices $u$ and $v$ are said to be adjacent. Otherwise, two vertices $u$ and $v$ are said to be nonadjacent. We use $E_{v}$ to denote the set of edges containing $v$ of $G.$ The degree of a vertex $v$ in $G$ is $d_{v} = |E_{v}|.$ In addition, we use $R(v)$ to denote the multiset of edge types in $E_{v}$ for a vertex $v.$

Let $G = (V, E)$  be a general hypergraph. The set $R=\{|e|\colon\ e \in E\}$ is called the set of edge types of $G.$ Then $G$ is also called an $R$-graph. For a set $S$ and a positive integer $i$, let $\binom{S}{i}=\{T\subseteq S \colon\ |T|=i\}.$ A complete $R$-graph on $n$ vertices is an $R$-graph with vertex set $[n]$ and edge set $\cup_{i\in R}\binom{[n]}{i},$ where $[n]=\{1,2,\ldots,n\}$. A clique of a $R$-graph $G$ is a complete $R$-subgraph of $G$. A maximal clique is a clique that cannot extended to a larger clique, and a maximal clique is a clique which covers as many vertices as possible. The clique number $\omega(G)$ of a hypergraph $G$ is the number of vertices in a maximum clique of $G$. If $E(G)=\varnothing$, then we define $\omega(G)=1.$

In 2005, the concept of tensor eigenvalues and the spectra of tensors were independently
introduced by Qi \cite{Q1} and Lim \cite{L}. A tensor $\mathcal{A}=(a_{i_{1}i_{2}\ldots i_{k}})$ with order $k$ and dimension $n$ is a multidimensional array, where $a_{i_{1}i_{2}\ldots i_{k}}\in \mathbb{C}$ and $1\leq i_{1},i_{2},\ldots,i_{k}\leq n.$ For an $n$-dimension complex vector $x=(x_{1}, x_{2}, \ldots, x_{n})^{T}$, $\mathcal{A} x^{k-1}$ is defined as an $n$-dimension complex vector whose $i$-th component is the following
$$(\mathcal{A}x^{k-1})_{i}=\sum\limits^{n}_{i_{2},\ldots,i_{k}=1}a_{ii_{2}\ldots i_{k}}x_{i_{2}}\cdots x_{i_{k}}, \mbox ~{for~any}~ i\in[n].$$

Let $x^{[k-1]}=(x_{1}^{k-1}, x_{2}^{k-1}, \ldots, x_{n}^{k-1})^{T}\in \mathbb{C}^{n}$. Then a number $\lambda\in \mathbb{C}$ is called an eigenvalue of the tensor $\mathcal{A}$ if there exists a nonzero vector $x\in \mathbb{C}^{n}$ such that $$\mathcal{A} x^{k-1}=\lambda x^{[k-1]},$$
and in this case, $x$ is called an eigenvector of $\mathcal{A}$ corresponding to the eigenvalue $\lambda.$ The spectral radius of $\mathcal{A}$ is $\rho(\mathcal{A}) = \max\{| \lambda | \colon\ \lambda \mbox{~is~an~eigenvalue~of~} \mathcal{A}\}.$ In particular, if $x$ is real, then $\lambda$ is called an $H$-eigenvalue. If $x\in\mathbb{R}^{n}_{+}$, then $\lambda$ is called an $H^{+}$-eigenvalue. If $x\in \mathbb{R}^{n}_{++},$ then $\lambda$ is called an $H^{++}$-eigenvalue \cite{Q2}.

\noindent\begin{definition}\label{de:c2}(\cite{BCM})
Let $G=(V,E)$ be a general hypergraph with $n$ vertices, $m$ edges and $rank(G)=k.$ The adjacency tensor $\mathcal{A}(G)$ of $G$ is  $$\mathcal{A}(G)=(a_{i_1,i_2,\ldots,i_k}),~~ 1\leq i_1,i_2,\ldots,i_k\leq n.$$
For all edges $e=\{v_{l_1},v_{l_2},\ldots,v_{l_s}\}\in E$ of cardinality $s\leq k,$
\begin{align*}
a_{i_1,i_2,\ldots,i_k}=\frac{s}{\alpha(s)},  \mbox ~~{where}~ \alpha(s)=\sum_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}}\frac{k!}{k_1!k_2!\cdots k_s!},
\end{align*}
and $i_1,i_2,\ldots,i_k$ are chosen in all possible ways from $L=\{l_1,\ldots,l_s\}$ with at least once for each element of the set $L,$ while each $k_{i}$ represents the times that $l_{i}$ appears in $\{ i_1,i_2,\ldots,i_k\}$. The other positions of the tensor are zero.
\end{definition}

The eigenvalues of a general hypergraph $G$ refer to the eigenvalues of the adjacency tensor of $G.$ And we use $\rho(G)$ to denote the spectral radius of a general hypergraph $G.$

Let $G$ be a general hypergraph with $n$ vertices and $rank(G)=k$, and $x=(x_{1}, x_{2}, \ldots, x_{n})^{T}$ be a vector. For an edge $e=\{i_{1},i_2,\ldots,i_s\}\in E,$ we have

\begin{align*}
x^{T}(\mathcal{A}(e)x^{k-1})=\frac{s}{\alpha(s)}\sum_{\substack{ k_1+\cdots+k_s=k,\\k_1,\ldots,k_s\geq 1}}\frac{k!}{k_1!k_2!\cdots k_s!}x_{i_1}^{k_1}x_{i_2}^{k_2}\cdots x_{i_s}^{k_s}.
\end{align*}
And we also have

\begin{align*}
x^{T}(\mathcal{A}(G)x^{k-1})=\sum_{e\in E(G)}x^{T}(\mathcal{A}(e)x^{k-1}).
\end{align*}

Let $\mathcal{D}(G)$ be a degree diagonal tensor with order $k$ and dimensional $n$, and its diagonal element $d_{ii\ldots i}$ (or simply $d_{i}$) be the degree of a vertex $v_{i}$ of $G$, for all $i \in \{1,2, \ldots, n\}$. Then $\mathcal{Q}(G)=\mathcal{D}(G)+\mathcal{A}(G)$ is the signless Laplacian tensor of the hypergraph $G$. The signless Laplacian eigenvalues of a hypergraph $G$ refer to the eigenvalues of the signless Laplacian tensor of $G$. We use $q(G)$ to denote the signless Laplacian spectral radius of $G$.

Similarly, Let $G$ be a hypergraph with $n$ vertices and $rank(G)=k$, and $x=(x_{1}, x_{2}, \ldots, x_{n})^{T}$ be a vector. For an edge $e=\{i_{1},i_2,\ldots,i_s\}\in E,$ we have

\begin{align*}
x^{T}(\mathcal{Q}(e)x^{k-1})=\sum_{j=1}^{s}x_{i_{j}}^{k}+\frac{s}{\alpha(s)}\sum_{\substack{ k_1+\cdots+k_s=k,\\k_1,\ldots,k_s\geq 1}}\frac{k!}{k_1!k_2!\cdots k_s!}x_{i_1}^{k_1}x_{i_2}^{k_2}\cdots x_{i_s}^{k_s}.
\end{align*}
And we also have

\begin{align*}
x^{T}(\mathcal{Q}(G)x^{k-1})=\sum_{e\in E}x^{T}(\mathcal{Q}(e)x^{k-1}).
\end{align*}

In 1965, Motzkin and Straus \cite{MS} defined a homogeneous polynomial $L(G, x)$ of a graph $G.$

\noindent\begin{definition}\label{de:c3}(\cite{MS})
Let $G=(V,E)$ be a graph with $n$ vertices, and let $x=(x_{1}, x_{2}, \ldots, x_{n})$ be a nonnegative real vector such that $\sum_{i=1}^{n}x_{i}=1$. Then $$L(G,x)=\sum_{\{i,j\}\in E}x_{i}x_{j}.$$
\end{definition}

Moreover, Motzkin and Straus \cite{MS} established a remarkable connection between the clique
number and a homogeneous polynomial $L(G)=\max\limits_{x\in S}L(G,x)$ of a graph $G,$ where $S=\left\{x \colon\ x \in R^{n}_{+}, \sum_{i=1}^{n} x_{i}=1\right\}.$

In 2020, Hou et al. \cite{HCZ} first defined a homogeneous polynomial $L(G, x)$ for a general hypergraph $G$, which generalized the definition $L(G, x)$ from a graph $G$ to a general hypergraph $G.$

\noindent\begin{definition}\label{de:c4}(\cite{HCZ})
Let $G=(V,E)$ be a general hypergraph with $n$ vertices and $rank(G)=k,$ and let $x=(x_{1}, x_{2}, \ldots, x_{n})$ be a nonnegative real vector such that $\sum_{i=1}^{n}x_{i}=1$. Then $$L(G,x)=\sum_{e\in E}\frac{1}{\alpha(s)}\sum_{\substack{ k_1+\cdots+k_s=k,\\k_1,\ldots,k_s\geq 1}}\frac{k!}{k_1!k_2!\cdots k_s!}x_{i_1}^{k_1}x_{i_2}^{k_2}\cdots x_{i_s}^{k_s},$$
where $s$ is the number of vertices of $e$ and $\alpha(s)=\sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1 }}\frac{k!}{k_1!k_2!\cdots k_s!}.$
\end{definition}

Similarly, for a general hypergraph $G,$  $$L(G)=\max_{x\in S}L(G,x),$$ where $S=\left\{x \colon\ x \in R^{n}_{+}, \sum_{i=1}^{n} x_{i}=1\right\}.$

In 1986, Wilf \cite{W} gave a lower bound on the spectral radius of a graph in terms of clique number, which was inspired by a result of Motzkin and Straus \cite{MS}. In 2007, Bollob\'{a}s and Nikiforov \cite{BN} obtained a number of relations between the number of cliques of a graph $G$ and the spectral radius. In 2007, Lu et al. \cite{LLT} presented some lower and upper bounds for the independence number and the clique number involving the Laplacian eigenvalues of a graph $G.$ In 2008, Liu and Liu \cite{LL} obtained some lower and upper bounds for the independence number and the clique number involving the signless Laplacian eigenvalues of a graph $G.$ In 2009, Rota and Pelillo \cite{RP} gave some new upper and lower bounds on the clique number of graphs. In 2009, Nikiforov \cite{N} gave some new bounds for the clique and independence numbers of a graph in terms of its eigenvalues. In 2013, He et al. \cite{HJZ} gave the bounds on the signless Laplacian spectral radius of graphs in terms of clique number, which disprove the two conjectures on the signless Laplacian spectral radius in \cite{HL}.

Recently, spectral hypergraph theory develops rapidly. There are many work about the
spectral theory of hypergraphs \cite{DW, DWL, DWXL, LZM, XW1, XW2}. However, there are still few studies the relation bewteen the clique number and the (signless Laplacian) spectral radius of hypergraphs. In 2015, Xie and Qi \cite{XQ} mainly researched some inequality relations between the signless (Laplacian) $H$-eigenvalues and the clique (coclique) numbers of uniform hypergraphs. In 2020, Hou et al. \cite{HCZ} first defined a homogeneous polynomial for a general hypergraph, and gave a Motzkin-Straus type result for $\{k-1, k\}$-graphs. And they gave some lower and upper bounds on the spectral radius of $\{k-1, k\}$-graphs in terms of clique number. In this paper, motivated by \cite{HCZ} and \cite{XQ}, we mainly generalize the results of Hou et al. \cite{HCZ} form $\{k-1,k\}$-graphs to $R$-graphs.

This paper is organized as follows. In Section 2, some necessary lemmas and inequalities are given. In Section 3, we give a
bound on the signless Laplacian spectral radius of general hypergraphs in terms of clique number. We also present a relation between a homogeneous polynomial and the clique number of general hypergraphs. Moreover, we also obtain a lower and a upper bound on the spectral radius of general hypergraphs in terms of clique number, respectively.

\section{Preliminaries}
In this section, we mainly give some useful lemmas and inequalities.

\noindent\begin{lemma}\label{le:ch-0}(Maclaurin's inequality \cite{HLP})
Let $x_{1}, x_{2}, \ldots , x_{n}$ be positive real numbers, and $$S_{k}=\frac{\sum\limits_{1 \leq i_{1}<i_{2}<\cdots<i_{k} \leq n} x_{i_{1}} x_{i_{2}} \cdots x_{i_{k}}}{\binom{n}{k}}, \text { for } 1 \leq k \leq n.$$
Then $$S_{1} \geq \sqrt{S_{2}} \geq \sqrt[3]{S_{3}} \geq \cdots \geq \sqrt[n]{S_{n}},$$
and the equality holds if and only if $x_{1} = x_{2} = \cdots = x_{n}.$
\end{lemma}

By Lemma \ref{le:ch-0}, we have $S_{1} \geq \sqrt[k]{S_{k}}$ for any $k \in [n].$ That is, $$\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k} \leq n} x_{i_{1}} x_{i_{2}} \cdots x_{i_{k}} \leq\binom{n}{k}\left(\frac{x_{1}+x_{2}+\cdots+x_{n}}{n}\right)^{k}.$$

\noindent\begin{lemma}\label{le:ch-1}(\cite{ZLKB})
(1). Let $G$ be a general hypergraph, then $\rho(G)$ is an $H^{+}$-eigenvalue of $G.$

(2). If $G$ is connected, then $\rho(G)$ is the unique $H^{++}$-eigenvalue of $G,$ with the unique eigenvector $x \in R^{n}_{++},$ up to a positive scaling coefficient.
\end{lemma}

If $G$ is a connected hypergraph, by Lemma \ref{le:ch-1}, then there exists a unique positive eigenvector $x=(x_{1}, x_{2}, \ldots, x_{n})^{T}$ corresponding to $\rho(G)$ such that $\sum_{i=1}^{n}x_{i}^{k}=1$, where $x$ is called the principal eigenvector of $\mathcal{A}(G)$. For a principal eigenvector $x$ of $\mathcal{A}(G),$ we have $$\rho(G)=x^{T} (\mathcal{A}(G) x^{k-1})=\sum_{e \in E(G)} x^{T} (\mathcal{A}(e) x^{k-1}).$$

Note that $\mathcal{Q}(G)$ is a weakly irreducible tensor. By the Perron-Frobenius theorem for nonnegative tensors \cite{FTP}, there exists a unique positive eigenvector $x=(x_{1}, x_{2}, \ldots, x_{n})^{T}$ corresponding to $q(G)$ such that $\sum_{i=1}^{n}x_{i}^{k}=1$, where $x$ is called the principal eigenvector of $\mathcal{Q}(G)$. Similarly, for a principal eigenvector $x$ of $\mathcal{Q}(G),$ we have $$q(G)=x^{T} (\mathcal{Q}(G) x^{k-1})=\sum_{e \in E(G)} x^{T} (\mathcal{Q}(e) x^{k-1}).$$

\noindent\begin{lemma}\label{le:ch-2}(\cite{HCZ})
Let $G$ be an $R$-graph with clique number $\omega.$ Then
$$\rho(G) \geq \sum_{s \in R}\binom{\omega-1}{s-1},$$
and the equality holds if and only if $G$ is a complete $R$-graph.
\end{lemma}

\section{ Main results}

The following Theorem \ref{th:c1} presents a lower bound on the signless Laplacian spectral radius of a general hypergraph in terms of clique number. This result generalizes Theorem 3.1 of Xie and Qi \cite{XQ} from uniform hypergraphs to general hypergraphs.

\noindent\begin{theorem}\label{th:c1}
Let $G$ be an $R$-graph with clique number $\omega.$ Then $$q(G) \geq 2\sum_{s \in R}\binom{\omega-1}{s-1},$$
and the equality holds if and only if $G$ is a complete $R$-graph.
\end{theorem}

\noindent\textbf{Proof.}
Assume that $G'$ is the maximum complete $R$-subgraph of $G$ with $\omega$ vertices and
$rank(G) = k.$ Let $x$ be a vector such that $x_{i}=\frac{1}{\sqrt[k]{\omega}}$ for $i \in V(G^{\prime})$ and $x_{i}=0$ for otherwise. It is obvious that $\|x\|_{k}^{k}=1.$  By the definition of the signless Laplacian spectral radius of general hypergraphs, we have

\begin{align}\label{eq:1}
q(G) & \geq x^{T} (\mathcal{Q(G)} x^{k-1}) \cr
& =\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}}(x_{i_{1}}^{k}+x_{i_{2}}^{k}+\cdots+x_{i_{s}}^{k}+\frac{s}{\alpha(s)} \sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1}! k_{2} ! \cdots k_{s}} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}}) \cr& =\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}}(\frac{1}{\omega}+\frac{1}{\omega}+\cdots+\frac{1}{\omega}+\frac{1}{\omega} \frac{s}{\alpha(s)} \sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !}) \cr
& =\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}}(\frac{s}{\omega}+\frac{1}{\omega} \frac{s}{\alpha(s)} \alpha(s)) \cr
& =\sum\limits_{s \in R} \sum\limits_{e \in E\left(G^{\prime}\right) \atop |e|=s} \frac{2 s}{\omega}=\sum\limits_{s \in R} \frac{2 s}{\omega}\binom{\omega}{s} \cr
&=2\sum\limits_{s \in R}\binom{\omega-1}{s-1}.
\end{align}

If $q(G)=2 \sum\limits_{s \in R}\binom{\omega-1}{s-1},$  by Inequality (\ref{eq:1}), then we have $q(G)= x^{T} \mathcal{Q}(G) x.$ Since Lemma \ref{le:ch-1} holds for $q(G)$ of a general hypergraph $G$, we know
$x$ is a positive vector. Therefore, we have $\omega = n.$ That is, $G$ is a complete $R$-graph.

If $G$ is a complete $R$-graph, then we have $\omega = n.$ By the definition of the signless Laplacian spectral radius of general hypergraphs, we have $q(G)=2 \sum\limits_{s \in R}\binom{\omega-1}{s-1}.$ \hfill$\square$

If $G$ be a $k$-uniform hypergraph with clique number $\omega$, by Lemma \ref{le:ch-2}, then we have $\rho(G)\geq \binom{\omega-1}{k-1}.$ Therefore, we only consider $R$-graphs such that $|R|>1$ in the following Theorem \ref{th:c2}.

\noindent\begin{theorem}\label{th:c2}
Let $G$ be an $R$-graph with clique number $\omega$, $rank(G)=k$ and $corank(G)=c.$ Then $$\rho(G) \geq \frac{(\omega-c+1)^{c-1}(\omega-k+1)^{k-c}}{(k-1) !}+\frac{(\omega-c+1)^{c-1}}{(c-1) !}.$$
\end{theorem}

\noindent\textbf{Proof.} By Lemma \ref{le:ch-2}, we have
\begin{equation*}\begin{aligned}
\rho(G) & \geq \sum_{s \in R}\binom{\omega-1}{s-1}\\
& \geq\binom{\omega-1}{k-1}+\binom{\omega-1}{c-1} \\
&=\frac{(\omega-1) !}{(k-1) !(\omega-k) !}+\binom{\omega-1}{c-1}\\
&=\frac{(\omega-1) !}{(c-1) !(\omega-c) !} \frac{(c-1) !(\omega-c) !}{(k-1) !(\omega-k) !}+\binom{\omega-1}{c-1} \\
&=\binom{\omega-1}{c-1} \frac{(c-1) !(\omega-c) !}{(k-1) !(\omega-k) !}+\binom{\omega-1}{c-1} \\& \geq \frac{(\omega-c+1)^{c-1}}{(c-1) !} \frac{(c-1) !(\omega-c) !}{(k-1) !(\omega-k) !}+\frac{(\omega-c+1)^{c-1}}{(c-1) !} \\
&=\frac{(\omega-c+1)^{c-1}}{(k-1) !}(\omega-c) \cdots(\omega-k+1)+\frac{(\omega-c+1)^{c-1}}{(c-1) !} \\
& \geq \frac{(\omega-c+1)^{c-1}(\omega-k+1)^{k-c}}{(k-1) !}+\frac{(\omega-c+1)^{c-1}}{(c-1) !},
\end{aligned}\end{equation*}
where $$\binom{\omega-1}{c-1}=\frac{(\omega-1) !}{(c-1) !(\omega-c) !}=\frac{(\omega-1) \cdots(\omega-c+1)}{(c-1) !} \geq \frac{(\omega-c+1)^{c-1}}{(c-1) !}.$$
\hfill$\square$

{\bf Remark.} If $c=k-1>1$, by Theorem \ref{th:c2}, then we have

\begin{equation*}
\begin{aligned}
\rho(G) & \geq \frac{(\omega-k+2)^{k-2}(\omega-k+1)}{(k-1) !}+\frac{(\omega-k+2)^{k-2}}{(k-2) !} \\
&=\frac{(\omega-k+2)^{k-2}}{(k-2) !}\left(\frac{\omega-k+1}{k-1}+1\right)\\
&=\omega \frac{(\omega-k+2)^{k-2}}{(k-1) !}\\
&>\frac{(\omega-k+2)^{k-1}}{(k-1) !}.
\end{aligned}
\end{equation*}

Thus, we know Theorem \ref{th:c2} improves Theorem 3.2 of Hou et al. \cite{HCZ}.

The following Theorem \ref{th:c3} generalizes Theorem 3.3 of Hou et al. \cite{HCZ} from $\{k-1, k\}$-graphs to $R$-graphs.

\noindent\begin{theorem}\label{th:c3}
Let $G=(V,E)$ be an $R$-graph with clique number $\omega$ and $rank(G)=k.$  If $G$ is a complete $R$-graph, or if there exists two nonadjacent vertices $v$ and $v'$ in $G$ such that $R(v)=R(v'),$ then $$L(G)=\sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s}.$$
\end{theorem}
\noindent\textbf{Proof.}
Assume that $G^{\prime }$ is the maximum complete $R$-subgraph of $G$ with $\omega$ vertices. Let $x$ be an $n$-dimensional vector such that $x_{i}=\frac{1}{\omega}$ for $i \in V(G^{\prime})$ and $x_{i}=0$ for otherwise. It is obvious that $\|x\|_{1}^{1}=1.$ By the definition of $L(G),$ we have

\begin{equation*}
\begin{aligned}
L(G) & \geq L(G, x) \\
&=\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}} \frac{1}{\alpha(s)} \sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}} \\
&=\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}} \frac{1}{\omega^{k}} \frac{1}{\alpha(s)}  \sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} \\
&=\sum\limits_{\substack{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E\left(G^{\prime}\right)\\s\in R}} \frac{1}{\omega^{k}} \frac{1}{\alpha(s)} \alpha(s) \\
&=\sum_{s \in R} \sum_{e \in E\left(G^{\prime}\right) \atop |e|=s} \frac{1}{\omega^{k}} \\
&=\sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s}.
\end{aligned}
\end{equation*}

In the following, we prove

\begin{align}\label{eq:2}
L(G) \leq \sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s}
\end{align}
by induction on $n$.

If $n \leq c-1,$ then the edge set of $G$ is empty. It is obvious that $L(G) \leq \sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s}$ holds. Assume that Inequality (\ref{eq:2}) holds for less than $n$ vertices. Without loss of generality, supposed that $n$ is sufficient large and $x$ is a vector such that $L(G)=L(G, x) .$ And $x_{1} \geq x_{2} \geq \cdots \geq x_{t}>x_{t+1}>\cdots>x_{n}=0.$

If $t<n,$ then we have $x_{n}=0 .$ We can obtain $G_{1}$ from $G$ by deleting the vertex $n$ and the edges containing the vertex $n$. Let $G_{1}$ be a general hypergraph with clique number $\omega_{1}$ and $\omega_{1} \leq \omega .$ By the induction hypothesis, we have

$$L(G)=L\left(G_{1}\right) \leq \sum_{s \in R}\frac{1}{\omega_{1}^{k}}\binom{\omega_{1}}{s} \leq \sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s},$$
where $\left(\frac{1}{\omega_{1}}\right)^{k}\binom{\omega_{1}}{s}$ is an monotonically increasing function of $\omega_{1}.$

If $t=n,$ then we have $x_{1} \geq x_{2} \geq \cdots \geq x_{n}>0$.

{\bf Case 1.} If $G$ is a complete $R$-graph, then we have $\omega=n .$ By the definition of $L(G)$ and triangle inequality, we have

\begin{align}\label{eq:3}
L(G) &=\max _{x \in S} L(G, x) \cr
&=\max _{x \in S} \sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G)\atop s\in R} \frac{1}{\alpha(s)} \sum_{ k_{1}+k_{2}+\cdots+k_{s}=k,\atop k_{1}, \ldots, k_{s} \geq 1} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}} \cr
& \leq \max _{x \in S} \sum_{e \in E(G)\atop |e|=c} \frac{1}{\alpha(c)} \sum_{k_{1}+k_{2}+\cdots+k_{c}=k,\atop k_{1}, \ldots, k_{c} \geq 1 } \frac{k !}{k_{1} ! k_{2} ! \cdots k_{c} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{c}}^{k_{c}} \cr
&+\cdots+\max _{x \in S} \sum_{e \in E(G) \atop|e|=k} x_{i_{1}} x_{i_{2}} \cdots x_{i_{k}}.
\end{align}

Without loss of generality, assume that there exist two integers $d \geq 1$ and $u \geq 0$ such taht $k=s d+u$ for any $s \in R .$ We know that

\begin{align}\label{eq:4}
\alpha(s) &=\sum_{k_{1}, \ldots, k_{s} \geq 1,\atop k_{1}+\cdots+k_{s}=k } \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} \cr
&=a_{1}\binom{s}{1}+a_{2}\binom{s}{2}+a_{3}\binom{s}{2}+\cdots+a_{b}\binom{s}{u} \cr
&=a_{1} \frac{s !}{(s-1) ! 1 !}+a_{2} \frac{s !}{(s-2) ! 2 !}+a_{3} \frac{s !}{(s-2) ! 2 !}+\cdots+a_{b} \frac{s !}{(s-u) ! u !} \cr
&=s !\left( \frac{a_{1}}{(s-1) ! 1 !}+ \frac{a_{2}}{(s-2) !2 !}+ \frac{a_{3}}{(s-2) !2 !}+\cdots+ \frac{a_{b}}{(s-u) ! u !}\right),
\end{align}
where $$a_{1}=\frac{k !}{(k-s+1) ! 1 ! \cdots 1 !},~~a_{2}=\left\{
\begin{array}{ll}
\frac{k !}{(k-s) ! 2 ! 1 ! \cdots 1 !},& \mbox {if}   ~k-s = 2,
\\
2! \frac{k !}{(k-s) ! 2 ! 1 ! \cdots 1 !},& \mbox {if}   ~k-s \neq 2,
\end{array}
\right. ,$$
and $$a_{3}=\left\{
\begin{array}{ll}
\frac{k !}{(k-s-1) ! 3 ! 1 ! \cdots 1 !},& \mbox {if}   ~k-s-1 = 3,
\\
2! \frac{k !}{(k-s-1) ! 3 ! 1 ! \cdots 1 !},& \mbox {if}   ~k-s-1 \neq 3,
\end{array}
\right. ,~~a_{b}=\frac{k !}{\underbrace{(d+1) ! \cdots(d+1) !}_{u~times} d ! \cdots d !}.$$

Since $G$ is a complete $R$-graph, we also have

\begin{align*}
&\sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop |e|=s}\sum_{k_{1}+k_{2}+\cdots+k_{s}=k,\atop k_{1}, \ldots, k_{s} \geq 1} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}}\\
&=\frac{k !}{(k-s+1) ! 1 ! \cdots 1 !} \sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G)} x_{i_{1}} x_{i_{2}} \cdots x_{i_{s}}\left(x_{i_{1}}^{k-s}+\cdots+x_{i_{s}}^{k-s}\right)\\
&+\frac{k !}{(k-s) ! 2 ! 1 ! \cdots 1 !} \sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G)} x_{i_{1}} x_{i_{2}} \cdots x_{i_{s}}\left(x_{i_{1}}^{k-s-1} x_{i_{2}}+\cdots+x_{i_{s-1}} x_{i_{s}}^{k-s-1}\right)+\cdots\\
&+\frac{k !}{(d+1) ! \cdots(d+1) ! d ! \cdots d !} \sum_{e=\left\{i_{1}, i_{2}, \dots, i_{s}\right\} \in E(G)} x_{i_{1}}^{d} x_{i_{2}}^{d} \cdots x_{i_{s}}^{d}\left(x_{i_{1}} \cdots x_{i_{u}}+\cdots+x_{i_{s-u+1}} \cdots x_{i_{s}}\right)\\
&=\frac{k !}{(k-s+1) ! }[x_{1}^{k-s+1}\sum_{e=\left\{1, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop 2\leq i_{2}< \cdots < i_{s}\leq n} x_{i_{2}}\cdots x_{i_{s}} +\cdots+x_{n}^{k-s+1}\sum_{e=\left\{i_{1}, i_{2}, \ldots, n\right\} \in E(G),\atop 1\leq i_{1}< \cdots < i_{s-1}\leq n-1} x_{i_{1}}\cdots x_{i_{s-1}} ]\\
&+\frac{k !}{(k-s) ! 2 !}[x_{1}^{k-s} x_{2}^{2}\sum_{e=\left\{1, 2, \ldots, i_{s}\right\} \in E(G),\atop 3\leq i_{3}< \cdots < i_{s}\leq n} x_{i_{3}}\cdots x_{i_{s}} +\cdots+x_{n-1}^{2} x_{n}^{k-s}\sum_{e=\left\{i_{1}, i_{2}, \ldots, n-1,n \right\} \in E(G),\atop 1\leq i_{1}< \cdots < i_{s-2}\leq n-2} x_{i_{2}}\cdots x_{i_{s}} ]\\
&+\cdots+\frac{k !}{(d+1) ! \cdots(d+1) ! d ! \cdots d !}[x_{1}^{d} \cdots x_{s}^{d}\sum_{e=\left\{i_{1}, i_{2}, \ldots,i_{u}, n-s+u+1,\ldots, n\right\} \in E(G), \atop 1\leq i_{1}< \cdots < i_{u}\leq n-s+u} x_{1}\cdots x_{u} \\ &+\cdots+ x_{1}^{d} \cdots x_{s}^{d}\sum_{e=\left\{1, 2, \ldots,s-u, i_{s-u+1},\ldots, i_{s}\right\} \in E(G),\atop s-u+1\leq i_{s-u+1}< \cdots < i_{s}\leq n} x_{i_{s-u+1}}\cdots x_{i_{s}}].
\end{align*}

Thus, by Maclaurin's inequality, we have
\begin{align*}
&\sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop |e|=s}\sum_{k_{1}+k_{2}+\cdots+k_{s}=k,\atop k_{1}, \ldots, k_{s} \geq 1} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}}\\
&\leq \frac{k !}{(k-s+1) ! } [x_{1}^{k-s+1}\left(\frac{\sum_{2 \leq i \leq n} x_{i}}{n-1}\right)^{s-1}\binom{n-1}{s-1}+\cdots+x_{n}^{k-s+1}\left(\frac{\sum_{1 \leq i \leq n-1} x_{i}}{n-1}\right)^{s-1}\binom{n-1}{s-1}]\\
&+\frac{k !}{(k-s) ! 2 ! }[x_{1}^{k-s} x_{2}^{2}\left(\frac{\sum_{3 \leq i \leq n} x_{i}}{n-2}\right)^{s-2}\binom{n-2}{s-2}+\cdots+x_{n-1}^{2} x_{n}^{k-s}\left(\frac{\sum_{1 \leq i \leq n-2} x_{i}}{n-2}\right)^{s-2}\binom{n-2}{s-2}]\\
&+\cdots+\frac{k !}{(d+1) ! \cdots(d+1) ! d ! \cdots d !}[x_{1}^{d} \cdots x_{s}^{d}\left(\frac{\sum_{1 \leq i \leq n-s+u} x_{i}}{n-s+u}\right)^{u}\binom{n-s+u}{u}\\
&+\cdots+x_{1}^{d} \cdots x_{s}^{d}\left(\frac{\sum_{s-u+1 \leq i \leq n} x_{i}}{n-s+u}\right)^{u}\binom{n-s+u}{u}].
\end{align*}
The equality holds if and only if $x_{1}=x_{2}=\cdots=x_{n}=\frac{1}{n} .$

For any $1\leq i \leq s$, we have $$\binom{n-i}{s-i}\binom{n}{i}=\binom{n}{s}\binom{s}{i}=\frac{n !}{(n-s) ! } \frac{1}{(s-i) !i !}.$$
Let $x_{1}=x_{2}=\cdots=x_{n}=\frac{1}{n}.$ Therefore, we have

\begin{align}\label{eq:5}
&\sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop |e|=s}\sum_{k_{1}+k_{2}+\cdots+k_{s}=k,\atop k_{1}, \ldots, k_{s} \geq 1} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}} \cr &
\leq a_{1}\left(\frac{1}{n}\right)^{k}\binom{n-1}{s-1}\binom{n}{1}+a_{2}\left(\frac{1}{n}\right)^{k}\binom{n-2}{s-2}\binom{n}{2}+\cdots+a_{b}\left(\frac{1}{n}\right)^{k}\binom{n-s+u}{u}\binom{n}{s-u} \cr &
=\left(\frac{1}{n}\right)^{k} \frac{n !}{(n-s) !}\left( \frac{a_{1}}{(s-1) ! 1 !}+ \frac{a_{2}}{(s-2) !2 !}+\cdots+ \frac{a_{b}}{u !(s-u) !}\right).
\end{align}

By (\ref{eq:4}) and (\ref{eq:5}), we have

\begin{align}\label{eq:6}
&\frac{1}{\alpha(s)} \sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G), \atop |e|=s} \sum_{k_{1}+k_{2}+\cdots+k_{s}=k,\atop k_{1}, \ldots, k_{s} \geq 1 }\frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}}\cr &\leq\left(\frac{1}{n}\right)^{k} \frac{n !}{(n-s) ! s !}=\left(\frac{1}{n}\right)^{k}\binom{n}{s}.
\end{align}

By (\ref{eq:3}) and (\ref{eq:6}), we have

$$L(G) \leq \sum_{s \in R}\left(\frac{1}{n}\right)^{k}\binom{n}{s}.$$

{\bf Case 2.} If there exist two nonadjacent vertices $v$ and $v'$ in $G$ such that $R(v)=R(v'),$ then we prove that there exists a vector $y$ such that $L(G, y) \geq L(G, x) .$ By the definition of $L(G),$ we denote $L^{v}(G)$ as

$$L^{v}(G)=\sum_{e \in E_{v},|e|=s\in R} \frac{1}{\alpha(s)} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} x_{i_{1}}^{k_{1}} x_{i_{2}}^{k_{2}} \cdots x_{i_{s}}^{k_{s}},$$
for any a vertex $v \in V(G).$

Let $y$ be a vector such that $y_{i}=x_{i}$ for $i \neq v, v',$ $y_{v}=x_{v}+x_{v'}$ and $y_{v'}=0 .$ And let $z$ be a vector such that $z_{i}=x_{i}$ for $i \neq v$ and $z_{v}=x_{v'} .$ It is obvious that $y, z \in R_{+}^{n} .$ For any positive integer $k,$ we have $\left(x_{v}+x_{v'}\right)^{k} \geq x_{v}^{k}+x_{v'}^{k} .$ Hence, we have $$L^{v}(G,y)\geq L^{v}(G,x)+L^{v}(G,z).$$ Moreover, without loss of generality, assume that $L^{v}(G,z)\geq  L^{v'}(G,x).$ Then we have $$L(G,y)- L(G,x)=L^{v}(G,y)- L^{v}(G,x)-L^{v'}(G,x)\geq L^{v}(G,z)-L^{v'}(G,x)\geq 0.$$

Thus, we can obtain $G_{1}$ from $G$ by deleting the vertex $v'$ and the edges containing $v'.$ Similarly, by the induction hypothesis, we have
$$L(G)=L(G,x)\leq L(G,y)\leq L(G_{1})\leq\sum_{s\in R}(\frac{1}{\omega})^{k}\binom{\omega}{s}.$$

Therefore, we have $$L(G)=\sum_{s \in R} \frac{1}{\omega^{k}}\binom{\omega}{s},$$
where $G$ is a complete $R$-graph, or there exists two nonadjacent vertices $v$ and $v'$ in $G$ such that $R(v) = R(v').$
\hfill$\square$

The following Theorem \ref{th:c4} generalizes Theorem 3.4 of Hou et al. \cite{HCZ} from $\{k-1, k\}$-graphs to $R$-graphs.

\noindent\begin{theorem}\label{th:c4}
Let $G = (V, E)$ be an $R$-graph with clique number $\omega$ and $rank(G) = k.$ If
$G$ is a complete $R$-graph, or if there exists two nonadjacent vertices $v$ and $v'$ in $G$ such that $R(v) = R(v'),$ then $$\rho(G)\leq\sum_{s\in R}k(\frac{U}{\omega})^{k}\binom{\omega}{s},$$
where $U$ is the sum of entries of the principal eigenvector.
\end{theorem}

\noindent\textbf{Proof.}
Let $x$ be the principal eigenvector of $G$ with $n$ vertices, and $y_{i}=\frac{x_{i}}{U},$ for $1\leq i \leq n.$ It is obvious that $\sum_{i=1}^{n}y_{i}=1.$  By Theorem \ref{th:c3} and the definition of $\rho(G),$ we have

\begin{align*}
\frac{\rho(G)}{U^{k}}&=\sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop s\in R}\frac{s}{\alpha(s)}\sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} y_{i_{1}}^{k_{1}} y_{i_{2}}^{k_{2}} \cdots y_{i_{s}}^{k_{s}}\cr
&\leq\sum_{e=\left\{i_{1}, i_{2}, \ldots, i_{s}\right\} \in E(G),\atop s\in R}\frac{k}{\alpha(s)}\sum\limits_{\substack{k_1+\cdots+k_s=k,\\ k_1,\ldots,k_s\geq 1}} \frac{k !}{k_{1} ! k_{2} ! \cdots k_{s} !} y_{i_{1}}^{k_{1}} y_{i_{2}}^{k_{2}} \cdots y_{i_{s}}^{k_{s}}\cr
&\leq \sum_{s\in R}k\frac{1}{\omega^{k}}\binom{\omega}{s}.
\end{align*}

Therefore, we have
$$\rho(G)\leq\sum_{s\in R}k(\frac{U}{\omega})^{k}\binom{\omega}{s}.$$
\hfill$\square$

\end{document}